Two dice are rolled. Find the probability of getting 5 on either die or the sum of both die is 5. I got 5/12 is that correct?
Answer by ewatrrr(24785)


Hi
Two dice are rolled:
P(getting 5 on either)= P(one being a 5 'and' one not being a 5)= 1/6*5/6
P(getting 5 on either 'or' sum of 5) = 1/6*5/6 + 4/36 = 9/36 = 1/4
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12